| (c) uti - cute = 0, a (t, 0) = , we (t, 0) = 0.

the question asks to determine the general solution
Answer #1

Ut - c²usse to is wave equation U60,0) = f(a) 471x) = g(0) By DAlembert is solution a+ct (6) ts ucx, t) = { [ f(x+et) +6Cx-ca r x² + 3 (²+2 7 + <IL 2xC xéte242 +2x6+ + g*+6°42-23/t - Je442242]++ COC+c+)?Q=C4}} = $2x [ **+36°42 ] + H { **42 +2264 – x